OpenStudy (anonymous):

Place the four steps of the elimination method in the proper order. Solve the resulting one-variable equation. Find the variable with equal or opposite coefficients. If equal, subtract the equations. If opposite, add the equations. Substitute the known value into the other equation and solve for the remaining variable.

OpenStudy (anonymous):

Okay, first, you subtract the equations if they are equal. As in: x + y + 2 = 3 2x + 2y + 2 = 8

OpenStudy (anonymous):

But if they are opposite, like this: x + y + 2 = 3 -2x - 2y - 2 = -4 You would add them. All you really have to do is combine like terms. So C = Step 1.

OpenStudy (anonymous):

Wait, B would be Step 1, my mistake. C is step 2.

OpenStudy (anonymous):

Then you solve for one of the variables, which is A

OpenStudy (anonymous):

And then you substitute that variable back into the equation, which is D.